Suppose your firm is considering investing in a project with the cash flows shown below, that the required rate of return on projects of this risk class is 11 percent, and that the maximum allowable payback and discounted payback statistics for your company are 3 and 3.5 years, respectively.

Time: 0 1 2 3 4 5 Cash flow: –$235,000 $65,800 $84,000 $141,000 $122,000 $81,200

Use the PI decision rule to evaluate this project. (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

NPV of Project = PV of Cash Flows - Investment = 65800/(1+11%)+84000/(1+11%)^2+141000/(1+11%)^3+122000/(1+11%)^4+81200/(1+11%)^5-235000 = 124106.98

PI index = (NPV+Investment)/Investment = 1+ 124106.98/235000 = 1.53

SInce PI is greater than 1 Project should be accepted.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well